OpenStudy (afloridagirl):

Arianna kicks a soccer ball off the ground and into the air with an initial velocity of 42 feet per second. Assume the starting height of the ball is 0 feet. Approximately how long does it take until the soccer ball hits the ground again? 1.3 secs 2.6 secs 16.0 secs 27.6 secs

OpenStudy (afloridagirl):

i cant decide between a or b

OpenStudy (afloridagirl):

@jim_thompson5910

OpenStudy (afloridagirl):

@agent0smith im sorry to tag you but can you help?

OpenStudy (afloridagirl):

pleasse help..

satellite73 (satellite73):

solve \[-16t^2+42t=0\] for \(t\)

OpenStudy (afloridagirl):

how do i do that?

satellite73 (satellite73):

factor out a \(-t\) first to get \[-t(16t-42)=0\]

satellite73 (satellite73):

ignore the first \(t\) because it is zero at 0 go with \[16t-42=0\]

satellite73 (satellite73):

so \[16t=42\] divide and you are done

OpenStudy (afloridagirl):

okay so 2.6?

satellite73 (satellite73):

yeah seems reasonable

OpenStudy (afloridagirl):

thank you!

satellite73 (satellite73):

yw

OpenStudy (afloridagirl):

it was wrong by the way... @satellite73